OpenStudy (pandagurl):

Maria will spin the arrow on the spinner 3 times. What is the probability that the arrow will stop on A, then B, then C? A. 3/256 B. 1/64 C. 1/12 D.3/4

OpenStudy (pandagurl):



OpenStudy (pandagurl):

I think C?

OpenStudy (pandagurl):

@K12awesomeness @reemii

OpenStudy (k12awesomeness):

it's b) 1/64

OpenStudy (pandagurl):

Can you explain please (The Unit test is gonna eat me alive if idk)

OpenStudy (k12awesomeness):

sure:)

OpenStudy (reemii):

B right, Because it's 3 times the same experiment. And in each experiment, P(A)=1/4, P(B)=1.4n P(C)=1/4 and P(D)=1/4. Therefore, "A then B then C" has a (1/4) * (1/4) * (1/4) probability of happening.

OpenStudy (k12awesomeness):

Well if those are the only 3 options of a segment, then there is a 1/3 chance of landing on A, then a 1/3 chance of landing on B, then a 1/3 chance of landing on C. Multiply the fractions together to get 1/27. However that isn't one of your answers. I'm thinking that the spinner has 4 sections, meaning 1/4 chance of A, then 1/4 chance of B, then 1/4 chance of C. Multiply these together to get 1/64, which is one of your answers. That's my input anyway

OpenStudy (k12awesomeness):

thats how I figured it out..

OpenStudy (pandagurl):

Thanks guys! :D